Circular No. 9047

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 K3 (BESHORE) E. Beshore reports his discovery of a comet with a diffuse nuclear

condensation in four co-added 30-s CCD images taken with the Mount Lemmon 1.5-m reflector (discovery observation tabulated below); co-added exposures show a faint 15′′ tail in p.a. ∼ 90^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, W. H. Ryan writes that CCD images obtained on May 28.25–28.27 UT with the Magdalena Ridge Observatory 2.4-m reflector also show a distinct tail in p.a. ≈ 90^{o}.

2009 UT α_{2000} δ_{2000} Mag. Observer
May 26.16484 11^{h}45^{m}40.^{s}39 +4^{o}22′33′′.6 20.5 Beshore

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2009-K62.

T = 2011 Jan. 9.266 TT  ω = 251.^{o}413
Ω = 0.032 }2000.0
q = 3.90156 AU  i = 146.680

DOUBLE QUASAR SDSS J153636.22+044127.0 R. Decarli and A. Treves, Università dell’Insubria, Como; R. Falomo,

Osservatorio Astronomico di Padova and INAF; M. Dotti, University of Michigan; M. Colpi, Università di Milano-Bicocca; and J. K. Kotilainen, Tuorla Observatory, University of Turku, report on a deep K-band image taken at the European Southern Observatory of the quasar SDSS J153636.22+044127.0 on Apr. 30.29 UT using the Very Large Telescope (+ HAWK-I). The spectrum of this radio-quiet QSO (z = 0.3893) exhibits two broad emission-line systems at slightly different redshifts (differing by 3300 km/s) in the Sloan Digital Sky Survey optical spectrum and was thus proposed as a candidate binary massive-black-hole system with sub-parsec separation (Boroson and Lauer 2009, Nature 458, 53). The new VLT image shows that the object is composed of two sources (with total magnitudes m_{K} = 14.1 and 15.8) at a separation of 1′′ (5.3 kpc). Both sources consist of a nucleus plus an extended emission; the host galaxies, with nuclei components subtracted, have total magnitudes m_{K} = 15.6 and 16.2. These results strongly suggest that SDSS J1536+0441 is a kpc-scale pair of quasars. This interpretation is consistent with the recent independent finding of two compact radio sources in the same source by Wrobel and

Laor (2009, http://lanl.arxiv.org/abs/0905.3566).

2009 May 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green